A firm produces 200 units of output at a total cost of $1,000. The firm's total fixed cost equals $200.The firm’s per-unit average variable cost is _____.

a.

​$4

b.

​$8

c.

$5

d.

​$10

e.

​$7

A game that demonstrates the basic problem confronting noncolluding oligopolists is known as _____.

a.

​a tit-for-strategy game

b.

​the prisoners’ dilemma

c.

​a cooperative game

d.

​a noncooperative game

e.

​the oligopolists’ dilemma

A linear demand curve for a good is:

a.

​always equal to one.

b.

​more price inelastic at higher price ranges.

c.

​always equal to zero.

d.

​more price elastic at higher price ranges.

e.

​more price elastic at lower price ranges.
Answer #1

Ans.1- (A)

Total cost = 1000

TFC = 200

So, total variable cost= 1000-200 = 800

So, average variable cost = 800/ 200= 4

Ans.2- (B)

A game that demonstrates the basic problem confronting noncolluding oligopolists is known as the prisoners dilemma

Ans.3- (D)

The elasticity along a linear demand curve is more elastic at higher price ranges and inelastic at lower price ranges.